OpenStudy (anonymous):

rewrite the following without an exponent (-7)^-1

OpenStudy (anonymous):

When your power is negative (-) get 1 and power will be pasetive (+) for example : 10^-1=1/10^1 7^-5=1/7^5 Now answer !:) :D

OpenStudy (anonymous):

1/7^1?

OpenStudy (anonymous):

Yeah !

OpenStudy (anonymous):

But its -7 and we have 1/-7^1

OpenStudy (anonymous):

oh okay, I forgot that lol. thanks!

OpenStudy (austinl):

Basically think of it like this, if you have a number to a negative exponent, you put that number under one to the positive exponent. \(\large{x^{-n}=\dfrac{1}{x^n}}\)

OpenStudy (anonymous):

okay thank you!

OpenStudy (anonymous):

can you help me with another one?

OpenStudy (austinl):

Who are you talking to, me or @E.ali ?

OpenStudy (anonymous):

(7/8)^-2 and it doesn't matter

OpenStudy (austinl):

\(\large{\dfrac{a}{b}^{-n}=\dfrac{b^n}{a^n}}\)

OpenStudy (anonymous):

so 7^2/8^2?

OpenStudy (austinl):

Nope, other way around :)

OpenStudy (anonymous):

but it says rewrite it with no exponents

OpenStudy (austinl):

Lol, what is \(8^2~and~7^2\)?

OpenStudy (anonymous):

64 and 49

OpenStudy (anonymous):

so its 64/49? that's the answer?

OpenStudy (austinl):

It is indeed!